# Unmasking Correlations in Nuclear Cross Sections with Graph Neural Networks

### v1

@@ Abstract
In this work, we explore the use of deep learning techniques to learn the relationships between nuclear cross-sections across the chart of isotopes. As a proof of principle, we focus on the neutron-induced reactions in the fast energy regime that are the most important in nuclear science and engineering. We use variational autoencoders (VAEs) and implicit neural representations (INRs) to build a learned feature representation space of nuclear cross sections and reduce the dimensionality of the problem. We then train graph neural networks (GNNs) on the resulting latent space to leverage the topological information encoded in the chart of isotopes and to capture the relationships between cross sections in different nuclei. We find that hypernetworks based on INRs significantly outperforms VAEs in encoding nuclear cross-sections. This superiority is attributed to INR’s ability to model complex, varying frequency details, which enables lower prediction errors when combined with GNNs. We also observe that GNN optimization is much more successful when performed in the latent space, whether using INRs or VAEs. However VAEs’ continuous representation also allows for direct GNN training in the original input space. We leverage these representational learning techniques and successfully predict cross sections for a $17\times 17$ block of nuclei with high accuracy and precision. These findings suggest that both representation encoding of cross-sections and the prediction task hold significant potential in augmenting nuclear theory models, e.g., providing reliable estimates of covariances of cross sections, including cross-material covariances.
@@ I Introduction
Measuring nuclear cross sections requires complex and expensive experimental facilities and is often impossible for short-lived radioactive nuclei that cannot be made as targets. While nuclear reaction theory has greatly improved over the past decades [12], its level of sophistication varies greatly depending on the nucleus and type of reaction. In very light nuclei, ab initio theory can successfully describe reactions at low energy in a fully quantum-mechanical framework [13]; in contrast, reactions of heavy nuclei with neutrons with just a few sub-eV energy already lead to the formation of a highly excited nucleus with an extremely complex structure that can only be described within a statistical framework [14, 15]. In practice, such statistical reaction theory relies on additional theoretical inputs describing nuclear properties such as the nuclear level density, the probability of capturing incident particles, the rate of decay through the emission of photons, neutrons, protons, etc. [16]. Today the best nuclear data libraries on the market largely rely on phenomenological models to determine such properties. These models are carefully calibrated to match realistic application requirements such as the criticality regimes of a nuclear reactor, at the price of a loss in generality and predictive power of the underlying theory [17, 4]. As a result, current models fail at consistently predicting the complex relationships between different types of cross sections in a single nucleus, or between cross sections in different nuclei across the chart of isotopes. This lack of predictive power is a significant impediment to efforts at designing new, cleaner or more efficient nuclear technologies. While research in fundamental reaction theory continues to progress, techniques originating from artificial intelligence and machine learning research offer an alternative viewpoint [18, 19, 20].
@@
The goal of this paper is to expose the intricate relationships between cross sections of different nuclei by leveraging such state-of-the-art deep learning models. To this end, we employ two complementary techniques in succession. As the first step, we use either a variational autoencoder (VAE) [24] or an implicit neural representation (INR) to uncover the latent space of nuclear cross sections and reduce the dimensionality of the problem. In the second step, we use GNNs operating on such latent spaces to predict how cross sections transform a nucleus $(Z,N)$ to a neighboring nucleus $(Z\pm 1,N\pm 1)$ in order to account for the implicit 2D geometric nature of the nuclear chart. The resulting network provides an efficient, consistent framework to predict unknown cross sections not too far from the training region.
@@
This paper is organized as follows: In Section II, we summarize the main aspects of our theoretical approach including the architecture of the network and the characteristics of the datasets used for training. Section III shows the results for different GNN architectures and encoder approaches. Section IV discusses our results in the context of related work on nuclear data before we summarize our findings in Section V.
II Methodology
@@ II.2.1 Variational Autoencoder
II.2.2 Implicit Neural Representation
@@
II.3 Graph Neural Networks
@@
$\mathcal{L}_{\rm MSE}=\|\chi(\mathcal{G})-\mathbf{y}_{z}\|_{2}\,,$ (4)
@@ III.1 Encoding Performance
Figure 4: Mean prediction error calculated using Eq. (8) of the GNN for each nucleus in the TENDL dataset. For better visualization, all plots are normalized between $\epsilon_{\rm min}$ and $0.1\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error over all plots. The three panels on the left side are based on VAE encoding while the three on the right on INR encoding. ARMA, Gated and GCN refer to the three types of GNNs considered in this work. The red square corresponds to the windowed test data: the GNN was not trained on it.
@@
where $\text{MSE}=\frac{1}{n}\sum^{n}_{i=1}\|\mathbf{y}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2}$, $\hat{\mathbf{x}}_{i}$ represents the reconstructed input generated by either the VAE or the INR, while $\mathbf{y}_{i}$ denotes the original cross-sectional inputs. Here, $n$ is the total number of testing samples. MAX represents the maximum possible value of the signal, which is equal to $1$. We evaluate each model using this metric on the entire test set. The VAE achieves a score of $43.67$, while the INR achieves a higher score of $47.20$. Overall, the performance of the INR is significantly better as this model is better able to precisely capture high-frequency information such as thresholds, for example the fact that the cross section for the $(n,2n)$ cross section is identically 0 unless the energy of the incident neutron exceeds the neutron separation energy.
III.2 GNN performance
The training of the GNN requires a partitioning of the TENDL dataset that is different from the one adopted in the training of the VAE and INR. Since the goal of the GNN, after training, is to make predictions of cross sections, possibly a few neutrons away from the training set, we divide the entire TENDL dataset into a training and test set based on a windowing function. The windowed test set is created using a rectangular window centered around a specific nucleus $(Z_{0},N_{0})$ in the chart of isotope and containing all nuclei with $Z_{0}-\Delta Z\leq Z_{0}\leq Z_{0}+\Delta Z$ and $N_{0}-\Delta N\leq N_{0}\leq N_{0}+\Delta N$. All results presented in this section were obtained for a test window centered on $(Z_{0},N_{0})=(42,60)$ with $\Delta Z=\Delta N=8$. This window contains a total of 289 test nuclei, 209 of which are included in the original TENDL dataset. Since that original dataset contains 2151 points on the 2D plane, this means that about 7.2% of the total data is allocated to the test set.
@@ III.2 GNN performance
where $\mathbf{\hat{x}}_{i}$ and $\mathbf{y}_{i}$ are the predicted node features and the ground truth in the original space, respectively, and $n$ is the actual number of nuclei in the set (either training or test). $(\mathbf{y}_{i}-\mathbf{\hat{x}}_{i})$ computes the point-wise difference and is a 256-dimension vector, which is then summed for each position $i$ in the test set. In the training phase $n=1942$ (the 2151 original nuclei of the TENDL dataset minus the 209 of the test window). During inference, the trained GNN is used to predict the node features in the test set and $n=209$. Since the model was only trained on the training data, there is no contamination bias from the test set on the GNN. The predicted cross-sections are then decoded using the chosen decoder $\mathcal{D}$, which is kept frozen to prevent any updates to the weights. We calculate the prediction error in the 256-dimension space using Eq. (8) and report it in Fig. 4 and Fig. 5.
In Fig. 4, the overall spread of the error is presented for all points in the dataset. We theorize that the traditional GCN architecture is too simple to estimate the structure of the TENDL grid. This is validated by the plots in Fig. 4 (bottom row), where the spread of the error in the GCN is quite large. On the other hand, the ARMA model shows excellent predictive performance in the test set for both the VAE and INR encoding. Interestingly, while the GatedConv model behaves as expected for the VAE encoding, we observe that for the case of the INR, there is a drop in performance. As we will discuss in Section IV, this is a consequence of the fact that the INR latent space is more restricted than the one from the VAE.
Since the GNN seeks to estimate the latent manifold to learn the underlying relationships of the graph, a narrow distribution of the latent variables hinders the GNN from learning these relationships accurately. To better quantitatively measure the performance of each network, we list in Table 1 the average test PSNR using Eq. (7) over all points in the test set. For completeness, we also provide comparisons with and without interpolation-based data augmentation, denoted as DA in the table.
@@ IV Discussion
The results presented in Section III pose a few interesting questions, both from a data science and from a nuclear physics perspective, which we discuss in this section.
IV.1 Representation learning and structure of the latent space
Figure 6: Visualization of the VAE and INR hypernet latent spaces. Panel (a): projections of each $k$-dimensional VAE latent vector onto a two-dimensional plane spanned by the two largest components of a principal component analysis. Panel (b): same figure for the INR latent vector. Panel (c): t-distributed stochastic neighbor (t-SNE) embedding of the $k$-dimensional VAE latent vector. Panel (d): same figure for the INR latent vector. In each panel, the color indicates the difference between neutron numbers vs. proton numbers, i.e., $N-Z$.
@@ IV.1 Representation learning and structure of the latent space
First, we emphasize the importance of the training objective of Eq. (4) for successful GNN training. When using the GNN with the original cross sections as targets (i.e., $\mathcal{L}_{\rm GNN}=\|\mathcal{D}(\chi(\mathcal{G}))-\mathbf{y}\|_{2}$), where $\mathcal{D}$ represents the VAE or INR decoder, and $\mathbf{y}$ is the original cross-sections with 256 dimensions, we observe that the GNN prediction performance is degraded with INR-based encoding compared to VAE. Unlike VAEs, which promote continuous representations through regularization (e.g., standard normal Gaussian prior), the INR’s latent space is more tightly constrained by its regularization term, particularly $\mathcal{L}_{\rm latent}$. This restriction may suppress features that would be discriminative for downstream prediction tasks. Additionally, the INR’s interconnected decoder architecture, where a hypernetwork generates the INR’s weights, could amplify the impact of perturbations in INR weights during GNN training, further hindering prediction performance.
@@
In Fig. 6, we show a direct visualization of the latent spaces of both the VAE and INR models. In the top panel, we first perform a principal component analysis (PCA) of the $k$-dimensional latent vectors and project each of them onto the two-dimensional plane spanned by the two largest PCA components. We first notice that the samples’ variability scale for the latent space in the INR model is significantly smaller than for its VAE counterpart: the INR latent space is much more “compact” than the VAE one. To better understand these structural difference we plot in the bottom panel of the figure the t-distributed stochastic neighbor (t-SNE) embedding. It shows that the latent space from the INR hypernetwork has very distinct clusters, which most likely correspond to more unevenness in its latent space transition compared to the VAE counterparts. In comparison, the VAE latent space is much more uniformly distributed.
We postulate that the smooth characteristics of the VAE’s latent space is what enables the GNN to build meaningful continuous representations when trained on the original cross sections instead of their latent representations. This likely contributes to its more stable prediction model, with better prediction quality, as compared to INRs in this setting. To fully harness the potential of INRs, future research should explore integrating desired properties into their latent space, enhancing its structure for downstream tasks and expanding their applicability. Such modification could extend the full potential of INRs for a wider range of applications.
IV.2 Predictions of unknown cross sections

### v2

# Learning nuclear cross sections across the chart of nuclides with graph neural networks

@@ Abstract
In this work, we explore the use of deep learning techniques to learn how nuclear cross-sections change as we add or remove protons and neutrons. As a proof of principle, we focus on the neutron-induced reactions in the fast energy regime that are the most important in nuclear science and engineering. We use variational autoencoders (VAEs) and implicit neural representations (INRs) to build a learned feature representation space of nuclear cross sections and reduce the dimensionality of the problem. We then train graph neural networks (GNNs) on the resulting latent space to leverage the topological information encoded in the chart of isotopes and to capture the variations of cross sections with the number of particles in nuclei. We find that hypernetworks based on INRs significantly outperforms VAEs in encoding nuclear cross-sections. This superiority is attributed to INR’s ability to model complex, varying frequency details, which enables lower prediction errors when combined with GNNs. We also observe that GNN optimization is much more successful when performed in the latent space, whether using INRs or VAEs. However VAEs’ continuous representation also allows for direct GNN training in the original input space. We leverage these representational learning techniques and successfully predict cross sections for a $17\times 17$ block of nuclei with high accuracy and precision. Using clustering algorithms, we map groups of latent vectors into regions of the nuclear chart and show that VAEs and INRs can discover some of the neutron magic numbers. These findings suggest that deep-learning models based on representation encoding of cross-sections combined with graph neural networks hold significant potential in augmenting nuclear theory models, e.g., providing reliable estimates of covariances of cross sections, including cross-material covariances.
@@ I Introduction
Measuring nuclear cross sections requires complex and expensive experimental facilities and is often impossible for short-lived radioactive nuclei that cannot be made as targets. While nuclear reaction theory has greatly improved over the past decades [12], its level of sophistication varies greatly depending on the nucleus and type of reaction. In very light nuclei, ab initio theory can successfully describe reactions at low energy in a fully quantum-mechanical framework [13]; in contrast, reactions of heavy nuclei with neutrons with just a few sub-eV energy already lead to the formation of a highly excited nucleus with an extremely complex structure that can only be described within a statistical framework [14, 15]. In practice, such statistical reaction theory relies on additional theoretical inputs describing nuclear properties such as the nuclear level density, the probability of capturing incident particles, the rate of decay through the emission of photons, neutrons, protons, etc. [16]. Today the best nuclear data libraries on the market largely rely on phenomenological models to determine such properties. These models are carefully calibrated to match realistic application requirements such as the criticality regimes of a nuclear reactor, at the price of a loss in generality and predictive power of the underlying theory [17, 4]. As a result, current models fail at consistently predicting the complex relationships between different types of cross sections in a single nucleus, or between cross sections in different nuclei across the chart of isotopes. This lack of predictive power is a significant impediment to efforts at designing new, cleaner or more efficient nuclear technologies. While research in fundamental reaction theory continues to progress, techniques originating from artificial intelligence and machine learning research offer an alternative viewpoint [18, 19, 20, 21].
@@
The goal of this paper is to expose the intricate relationships between cross sections of different nuclei by leveraging such state-of-the-art deep learning models. To this end, we employ two complementary techniques in succession. As the first step, we use either a variational autoencoder (VAE) [25] or an implicit neural representation (INR) to uncover the latent space of nuclear cross sections and reduce the dimensionality of the problem. In the second step, we use GNNs operating on such latent spaces to predict how cross sections transform when adding a neutron or a proton to a nucleus $(Z,N)$ in order to account for the implicit 2D geometric nature of the nuclear chart. The resulting network provides an efficient, consistent framework to predict unknown cross sections not too far from the training region.
@@
This paper is organized as follows: Section II gives a broad overview of the challenges facing nuclear theory when computing nuclear data. In Section III, we summarize the main aspects of our theoretical approach including the architecture of the network and the characteristics of the datasets used for training. Section IV shows the results for different GNN architectures and encoder approaches. Section V discusses our results in the context of related work on nuclear data before we summarize our findings in Section VI.
II Theoretical models of nuclear data: Challenges
As mentioned in the introduction, nuclear reaction libraries are formed by the set of all known cross sections for all known nuclei. Most of this data is not simply a set of experimental data points but the result of an evaluation of all available data by subject matter experts. This evaluation is itself the outcome of a complex, multi-stage, iterative process that involves: (i) carefully analyzing available experimental results (vetting outliers, estimating uncertainties, etc.) (ii) adjusting model parameters of nuclear reaction codes such as TALYS [26] or EMPIRE [27] in such a way that the cross sections computed with such codes “match”, or are close to, experimental measurements (iii) processing the data (energy binning, heating of cross sections to room temperature values, etc.) (iv) verifying the nuclear reaction library through benchmarks, e.g., running a suite of codes simulating critical assemblies.
[…]
In practice, the physics of a nuclear reaction is instead described by a suite of models with varying degrees of sophistication and phenomenology. These models play an essential role in the evaluation process and can be divided into several broad categories: (i) optical models describe the interaction of an incoming particle $a$ with a target $A$ (ii) level density models characterize the number of available energy levels $\rho(E,J,\pi)$ at a given excitation energy $E$, angular momentum $J$ and parity $\pi$ (iii) gamma-strength functions $\gamma_{SF}(E)$ characterize the probabilities of emitting photons as a function of energy $E$ (iv) fission transmission coefficients quantify the probability for a (heavy) nucleus to undergo fission as a function of energy, angular momentum and parity (v) pre-equilibrium reaction models describe interaction mechanisms between projectile and target that involve specific nuclear excited states.
All these models are but different facets of the ideal quantum-mechanical evolution describe above. This implies that they should all be related to one another, at the very least indirectly. For example, let us assume one were to describe the structure of an atomic nucleus with a self-consistent mean-field model based on some effective nucleon-nucleon potential [28]. How this mean field is computed or parametrized should directly impact the optical potential (the real part of the optical potential does represent the nuclear mean field); the nuclear level density (through the excited states associated with the mean-field potential); the gamma-strength-function (which is typically related to functional derivatives of the mean field with respect to nuclear densities); fission transmission coefficients (which depend on how the mean-field potential changes with deformation), etc. Moreover, since the same effective nucleon-nucleon potential would be used to compute every single nucleus, nuclear properties would also be heavily correlated as a function of proton and neutron number.
Therefore, even if such a theoretical model still contains parameters fitted to experimental data – in this case: the parameters of the effective nucleon-nucleon force, all nuclear properties are directly related to that single set of parameters. In a sensitivity analysis, one would find that variations of the model parameters result in correlated variations of the model outputs, i.e., the nuclear data properties of interest. In current approaches based on separate models for each type of nuclear data, such correlated variations, which manifest relationships between different data, are impossible to capture with nuclear theory alone.
Instead, each nuclear model is treated as a separate black box independently parametrized by a handful of parameters. In a typical evaluation process, these parameters are adjusted in such a way that nuclear cross sections match experimental measurements and the resulting benchmarks are satisfied [3, 4, 29, 2]. These fits will often take the form of a functional dependency of model parameters in terms of proton and neutron number (or atomic mass number $A$), excitation energy or angular momentum. The nuclear reaction codes mentioned above typically implement a few different variants of such systematics. By construction, these systematics become highly questionable in nuclei or at energies that are beyond the range of their fit, which directly impacts the reliability of nuclear data libraries in these regions.
By employing a collection of independent nuclear models locally calibrated to experimental data, it thus becomes nearly impossible to (1) make reliable predictions outside the range of currently available measurements (2) capture the relationships (or correlations) between nuclear data within the same nucleus, or between different nuclei. In this paper, we will show that deep-learning techniques provide an alternative solution to point (2) .
@@ III.2.1 Variational Autoencoder
III.2.2 Implicit Neural Representation
@@
III.3 Graph Neural Networks
@@
$\mathcal{L}_{\rm GNN}=\|\chi(\mathcal{G})-\mathbf{y}_{z}\|_{2}\,,$ (4)
@@ IV.1 Encoding Performance
Figure 4: Mean prediction errors calculated using Eq. (8) of the GNN for each nucleus in the TENDL dataset. For better visualization, all plots are normalized between $\epsilon_{\rm min}$ and $0.1\times\epsilon_{\rm max}$, where $\epsilon_{\rm min}$ ($\epsilon_{\rm max}$) is the minimum (maximum) error over all plots. The three panels on the left side are based on VAE encoding while the three on the right on INR encoding. ARMA, Gated and GCN refer to the three types of GNNs considered in this work. The red square corresponds to the windowed test data: the GNN was not trained on it.
@@
where $\text{MSE}=\frac{1}{n}\sum^{n}_{i=1}\|\mathbf{y}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2}$, $\hat{\mathbf{x}}_{i}$ represents the reconstructed input generated by either the VAE or the INR, while $\mathbf{y}_{i}$ denotes the original cross-sectional inputs. Here, $n$ is the total number of testing samples. MAX represents the maximum possible value of the signal, which is equal to $1$. We evaluate each model using this metric on the entire test set. The VAE achieves a score of $43.67$, while the INR achieves a higher score of $47.20$. Overall, the performance of the INR is significantly better as this model is better able to precisely capture high-frequency information such as thresholds, for example the fact that the cross section for the $(n,2n)$ cross section is identically 0 unless the energy of the incident neutron exceeds the neutron separation energy. For completeness, Appendix VII.1 shows the evolution of the train and test losses as well of the train and test PSNR as a function of epochs.
IV.2 GNN performance
The training of the GNN requires a partitioning of the TENDL dataset that is different from the one adopted in the training of the VAE and INR. Since the goal of the GNN, after training, is to make predictions of cross sections, possibly a few neutrons away from the training set, we divide the entire TENDL dataset into a training and test set based on a windowing function. The windowed test set is created using a rectangular window centered around a specific nucleus $(Z_{0},N_{0})$ in the chart of isotope and containing all nuclei with $Z_{0}-\Delta Z\leq Z_{0}\leq Z_{0}+\Delta Z$ and $N_{0}-\Delta N\leq N_{0}\leq N_{0}+\Delta N$. All results presented in this section were obtained for a test window centered on $(Z_{0},N_{0})=(42,60)$ with $\Delta Z=\Delta N=8$. This window contains a total of 289 test nuclei, 209 of which are included in the original TENDL dataset. Since that original dataset contains 2151 points on the 2D plane, this means that about 7.2% of the total data is allocated to the test set. Appendix VII.2 gives results obtained with different window center locations and different window sizes.
@@ IV.2 GNN performance
where $\mathbf{\hat{x}}_{i}$ and $\mathbf{y}_{i}$ are the predicted node features and the ground truth in the original space, respectively, and $n$ is the actual number of nuclei in the set (either training or test). The quantity $(\mathbf{y}_{i}-\mathbf{\hat{x}}_{i})$ computes the point-wise difference and is a 256-dimension vector, which is then summed for each position $i$ in the test set. In the training phase $n=1942$ (the 2151 original nuclei of the TENDL dataset minus the 209 of the test window). During inference, the trained GNN is used to predict the node features in the test set and $n=209$. Since the model was only trained on the training data, there is no contamination bias from the test set on the GNN. The predicted cross-sections are then decoded using the chosen decoder $\mathcal{D}$, which is kept frozen to prevent any updates to the weights. We calculate the prediction error in the 256-dimension space using Eq. (8) and report it in Fig. 4 and Fig. 5.
In Fig. 4, the overall spread of the error is presented for all points in the dataset. We theorize that the traditional GCN architecture is too simple to estimate the structure of the TENDL grid. This is validated by the plots in Fig. 4 (bottom row), where the spread of the error in the GCN is quite large. On the other hand, the ARMA model shows excellent predictive performance in the test set for both the VAE and INR encoding. Interestingly, while the GatedConv model behaves as expected for the VAE encoding, we observe that for the case of the INR, there is a drop in performance. As we will discuss in Section V, this is a consequence of the fact that the INR latent space is more restricted than the one from the VAE.
Since the GNN seeks to estimate the latent manifold to learn the underlying relationships of the graph, a narrow distribution of the latent variables hinders the GNN from learning these relationships accurately. To better quantitatively measure the performance of each network, we list in Table 1 the average test PSNR using Eq. (7) over all points in the test set. We also provide comparisons with and without interpolation-based data augmentation, denoted as DA in the table.
@@ V Discussion
The results presented in Section IV pose a few interesting questions, both from a data science and from a nuclear physics perspective, which we discuss in this section. In particular, we focus on a brief analysis of the structure of the latent space to analyze if specific patterns can be matched to nuclear properties. We also outline how our framework lends itself to propagating uncertainties of cross sections throughout the nuclear chart.
V.1 Representation learning and structure of the latent space
@@
First, we emphasize the importance of the training objective of Eq. (4) for successful GNN training. When training the GNN with the original cross-sections as targets (i.e., $\mathcal{L}_{\rm GNN}=\|\mathcal{D}(\chi(\mathcal{G}))-\mathbf{y}\|_{2}$), where $\mathcal{D}$ represents the VAE or INR decoder, and $\mathbf{y}$ is the original cross-sections with 256 dimensions, we observe that the GNN prediction performance is degraded with INR-based encoding compared to VAE. Unlike VAEs, which promote continuous representations through regularization (e.g., standard normal Gaussian prior), the INR’s latent space is more tightly constrained by its regularization term, particularly $\mathcal{L}_{\rm latent}$. This restriction may suppress features that would be discriminative for downstream prediction tasks. Additionally, the INR’s interconnected decoder architecture, where a hypernetwork generates the INR’s weights, could amplify the impact of perturbations in INR weights during GNN training, further hindering prediction performance.
Figure 6: PCA visualization of the VAE and INR hypernet latent spaces. Panel (a): projections of each $k$-dimensional VAE latent vector onto a two-dimensional plane spanned by the two largest components of the PCA. Panel (b): same figure for the INR latent vector.
@@
In Fig. 6, we show a direct visualization of the latent spaces for the training dataset of both the VAE and INR models. In both panels, we perform a principal component analysis (PCA) of the $k$-dimensional latent vectors ($k=32$) and project each of them onto the two-dimensional plane spanned by the two largest PCA components. Panel (a) corresponds to the VAE latent space while panel (b) corresponds to the INR latent space. We recall that each point in such a figure corresponds to a given nucleus; its color is the value of the difference $N-Z$ of the neutron number $N$ and proton number $Z$ in the corresponding nucleus. The only noticeable difference is that the samples’ variability scale for the latent space in the INR model is significantly smaller than for its VAE counterpart: the INR hypernet latent space is much more “compact” than the VAE one.
Figure 7: t-SNE visualization of the VAE and INR hypernet latent spaces. Panel (a): two-dimensional t-SNE embedding of $k$-dimensional VAE latent vectors obtained for a perplexity value of 200. Panel (b): same figure for the INR latent vector.
To better understand the structural differences between the two encoding techniques, we compute the t-distributed stochastic neighbor (t-SNE) embedding [43]. Let us quickly recall the basics of the t-SNE method. Assume we have a set $\mathcal{P}$ of points $P$ in a $n_{p}$-dimensional space, and characterize the “similarity” of two points $P$ and $P^{\prime}$ in $\mathcal{P}$ by some joint probability distribution function $p$. Let us now assume that $\mathcal{P}$ can be mapped onto a set $\mathcal{Q}$ of points $Q$ in a $n_{q}$-dimensional space ($n_{q}\ll n_{p}$) that is characterized by another probability distribution $q$. The t-SNE technique minimizes the Kullback-Leibler divergence $\mathcal{L}_{PQ}$ between the joint probabilities $p$ in the high- dimensional space and the joint probabilities $q$ in the low-dimensional space. In simple terms: two points that are “neighbors” in the high-dimensional space should also be neighbors in the low-dimensional one.
Figure 7 shows the two-dimensional representation of this t-SNE embedding for both the VAE and INR encodings as obtained with the scikit-learn implementation. Like Fig. 6, the color of a point is the difference $N-Z$ of the neutron number $N$ and proton number $Z$ in the corresponding nucleus. In practice, t-SNE algorithms require setting a few hyperparameters, most notably one called “perplexity” ${\tt Perp}$ which can be interpreted as “a smooth measure of the effective number of neighbors” in the high-dimensional space [43]. Although traditional implementations of t-SNE suggest values of the perplexity between 5 and 50, this is highly dependent on the dataset and its dimensionality. In our case, the dataset has 1936 points – training nuclei represent 90% of the whole TENDL dataset of 2151 nuclei – and is encoded in a latent space of dimension $k=32$. We found that in order to converge the Kullback-Leibler loss $\mathcal{L}_{PQ}$, the perplexity parameter had to be significantly larger: figure 7 was obtained for ${\tt Perp}=200$. Note that the spatial arrangement of points and the variability scale in each latent dimension are very dependent on the value of ${\tt Perp}$. In particular, small values of ${\tt Perp}$ (corresponding to large values of $\mathcal{L}_{PQ}$) lead to a lot of clustering and more compact latent spaces.
Figure 8: Location of each nucleus in the training set with color code based on which of 4 clusters the nucleus belong too. Black horizontal and vertical lines represent magic numbers. Black squares are stable nuclei.
To better analyze clustering structures, we used the scikit-learn implementation of the spectral clustering algorithm to identify potential clusters of points directly in the VAE or INR dataset. Out of all available clustering techniques, spectral clustering and the DBScan algorithms are the naive best choices to identify clusters based on existing documentation and literature. However, calibrating the maximum distance between any two points to be considered in the neighborhood of each other in DBscan requires a deeper understanding of the latent space than we currently have. In contrast, spectral clustering does not involve any prior knowledge of the latent space. We first determined a rough estimate of the optimal number $n_{\rm clus}$ of clusters based on the Silhouette method. In practice, we found $n_{\rm clus}=4$. Figure 8 shows the position in the nuclear chart of each nucleus from the INR encoding training dataset, color-coded by the cluster the nucleus belongs to. The most striking feature of this plot is the fact that nuclei are organized in diagonal bands loosely following the valley of stability rather than along isotopic, isotonic or isobaric sequences. In addition, some of the clusters seem to be related to neutron magic numbers (but, interestingly, not proton magic numbers), especially $N=82$ and $N=126$. Both observations are quite robust. In Appendix VII.3, we show that different clustering scenarios lead to similar qualitative results.
V.2 Predictions of unknown cross sections